OpenStudy (anonymous):

Could somebody help please? I will give a medal!(: (Hold on, i will upload the pic)

OpenStudy (anonymous):

Its just #5 & #7 I'm having trouble with.

OpenStudy (anonymous):

5 is tricky it looks like there is a repeat in the first coordinate, those two points whose first coordinate is something like \(-1.5\)

OpenStudy (anonymous):

so i would say "not a function" for that one as for #7 \[f(x)=-x+4\]so \[f(-2)=-(-2)+4=2+4=6\] etc

OpenStudy (anonymous):

How do you add all that. Im not sure on how to do f(-)= -(-2) part @satellite73

OpenStudy (anonymous):

would it just be that equation?

OpenStudy (amistre64):

f(x) is defined already by an equation when x = -2, we can then determine the value for f(-2)

OpenStudy (anonymous):

okay thanks!